From GlueXWiki
Jump to: navigation, search

January 7, 2016 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. NIM Papers
  2. Spring 2016 Run Planning Meeting
  3. CDC update (Beni/Mike)
  4. FDC update (Lubomir)
  5. F125 Update (Naomi)
  6. Electronics (Fernando, Chris, Nick)
  7. Other


Participants: Curtis, Naomi, Mike (CMU), Sean (NU), Fernando, Eugene, Luke, Chris, Nick, Simon, Beni, and Lubomir (JLab).

Spring 2016 Run Planning Meeting

- In preparation for the coming run need to perform complete tests of the detectors: should start taking cosmic data early next week. Expect to have physics data taking at least at the end of the run (once the calorimeters are calibrated).

- FDC was turned on few days ago. One positive HV channel (second package, most downstream cell, side HV sector) draws huge current, can be operated at 200V below nominal. The problem started in early December. Most likely it's outside of the gas volume since the filed HV is not affected. It has been checked that it is not the HV module. Fixing will require access to the HV connection (at the bottom of the second package) and it can't be done now. If needed (still 1% of the sensitive area) can be done during the summer. CDC has not been turned on yet.

F125 Update

- Naomi reported on the test of the f125 firmware status (see link above). She compared the emulated results using the raw data with the firmware quantities. After fixing with Cody several problems, only a few and very rare issues remain as listed in the link. The issue with the short events initially caused crashing of the analyzing software but was eventually fixed by David L.

- Eugene: analyzer shouldn't crash, has to have some checks of the buffer sizes. Fernando: important is to understand if the errors are random (most likely timing issues in the firmware), or related mostly to a certain hardware (hardware issues).

FDC update

- During the last run there was an attempt to set HV on the FDC without magnetic field. As explained in the entry above by Luke: many HV channels tripped when they switched to 10^-4 radiator at 50nA current at 2050V (compared to 2200V nominal) with 1uA trip limit. Interestingly half of the channels were not central, i.e. with HV sectors just next to the beam, but >20cm on the side.

- Results from the TRD test in the last run are very promising. The first two attached plots show the amplitude (proportional to dE/dx) vs drift time and horizontal coordinate for two radiator configurations. One can clearly sees the places with radiator: higher amplitudes especially for big drift time. This is also shown when comparing the amplitudes vs drift time with and without radiator (next two plots). The difference between the two (last plot) comes from the transition radiation photons and as a function of the drift time (roughly proportional to photon interaction point along the electron track) is consistent with Sergey's simulations.

NIM papers

- This was a reminder that it is now time to start writing NIM papers on the two detectors. The tex templates were prepared by Beni some time ago (see 11/19/2015 meeting minutes).